OpenStudy (anonymous):

6 5/8%...convert to decimal notation.must show work

OpenStudy (anonymous):

0.625?

OpenStudy (anonymous):

arent we suppsed to divide it by 100 also?

OpenStudy (anonymous):

you can just turn *5 into a decimal by moving the decimal. So you can do 65/.08

OpenStudy (anonymous):

0.06625?

OpenStudy (anonymous):

or 0.6625?

OpenStudy (aripotta):

that's what i got :/

OpenStudy (aripotta):

0.06625

OpenStudy (anonymous):

0.625?

OpenStudy (anonymous):

how did you get that tho?my thought was to move the decimal places over 2 times to the left becuase of the dividing it by 100

OpenStudy (anonymous):

my teacher said 0.625 was wrong and that i had to divide it by 100. but im not there anymore to ask

OpenStudy (anonymous):

i will go with my gut and accept 0.06625 becuase thats what i got too

OpenStudy (aripotta):

6 5/8% = 53/8% = 6.625% = 0.06625

OpenStudy (anonymous):

thanks everyone:)